Suppose the price of commodity X falls from N15 per kg to N10 per kg and the quantity demanded increases from 100 kg to 300 kg. Find the price elasticity of demand.


1
Expert's answer
2022-04-15T16:38:46-0400

Price Elasticity of Demand = percent change in quantity / percent change in price.

deltaN = 100 x (15-10)/10 = 50%

deltaP = 100 x (300-100) / 100 = 200%

Price Elasticity of Demand = 200 / 50 = 4